OpenStudy (anonymous):

Find a function that satisfies the given conditions: f(0) = f(1) =0; integral f(x)dx from 0 to 1 = 2; f(x) >=0 for all x f'(x) exists for all x

OpenStudy (zarkon):

start by looking at the function \[g(x)=x^2(1-x)^2\]

OpenStudy (zarkon):

you will gave to modify \(g\) a little to get the function \(f\) that you want.

OpenStudy (anonymous):

Zarkon, how did you arrive at that initial function, please?

OpenStudy (zarkon):

using the first conditions (f(0) = f(1) =0) i thought of g(x)=x(1-x) you also wanted the function to be non-negative for all x so I squared the function it is a pollynomial so it is differentiable. the only part to do is to make it integrate to 2

OpenStudy (zarkon):

so find a constant c such that \[\int_{0}^{1}cx^2(1-x)^2dx=2\]

OpenStudy (zarkon):

then you will have \(f\)

OpenStudy (anonymous):

Sweet! I'll try that. What I'm really trying to arrive at is a function which gives the minimal arc length for a curve which satisfies those conditions. Thanks.